# Self-MoE: Towards Compositional Large Language Models with Self-Specialized Experts

### v1

@@ Abstract
We present Self-MoE, an approach that transforms a monolithic LLM into a compositional, modular system of self-specialized experts, named MiXSE (MiXture of Self-specialized Experts). Our approach leverages self-specialization, which constructs expert modules using self-generated synthetic data, each equipped with a shared base LLM and incorporating self-optimized routing. This allows for dynamic and capability-specific handling of various target tasks, enhancing overall capabilities, without extensive human-labeled data and added parameters. Our empirical results reveal that specializing LLMs may exhibit potential trade-offs in performances on non-specialized tasks. On the other hand, our Self-MoE demonstrates substantial improvements over the base LLM across diverse benchmarks such as knowledge, reasoning, math, and coding. It also consistently outperforms other methods, including instance merging and weight merging, while offering better flexibility and interpretability by design with semantic experts and routing. Our findings highlight the critical role of modularity and the potential of self-improvement in achieving efficient, scalable, and adaptable systems. Our code will be released upon acceptance.
Self-MoE: Towards Compositional Large Language Models
with Self-Specialized Experts
Junmo Kangthanks: Work done during internship at MIT-IBM Watson AI Lab.1,2   Leonid Karlinsky2   Hongyin Luo3   Zhen Wang4,5 Jacob Hansen3   James Glass3   David Cox2   Rameswar Panda2 Rogerio Feris2   Alan Ritter1 1Georgia Institute of Technology   2MIT-IBM Watson AI Lab 3Massachusetts Institute of Technology  4UC San Diego  5MBZUAI junmo.kang@gatech.edu
@@ 1 Introduction
Figure 1: Concept of Self-MoE, illustrating the transformation from a monolithic LLM to a compositional system, MiXSE, without extensive resources and addition of significant parameters. The results showcase MiXSE’s improved capabilities over the base LLM (e.g., Gemma-7B) across all domains, unlike the knowledge-specialized LLM that compromises other capabilities.
Meanwhile, the increasing demand to handle domain-specific or expert-level tasks has highlighted the need for specialization of LLMs (Cheng et al., 2024; Ling et al., 2023; Feng et al., 2024). However, effective tuning often relies on high-quality, human-annotated data, which is costly and challenging to scale (Kang et al., 2023b), especially in specialized domains where expertise is scarce and valuable (Wu et al., 2023). Self-specialization (Kang et al., 2023a) offers a promising alternative, aligning models with self-generated synthetic data. While this technique has proven effective in cross-task generalization within a target expert domain, we posit that it may compromise performance in areas outside the target domain.
In this paper, we explore the following question: How can we build compositional LLMs that enjoy versatile expertise, while using minimal resources? We introduce Self-MoE (Figure 1), an approach that transforms a monolithic model into a compositional (Zaharia et al., 2024) system, called MiXSE (MiXture of Self-specialized Experts). This approach differs from prior MoE work using LoRA (Hu et al., 2022), which either relies on human-labeled data (Wu et al., 2024) or assumes the existence of trained modules (Huang et al., 2023; Muqeeth et al., 2024). Instead, our Self-MoE constructs individual lightweight expert modules from scratch using synthetic data, inspired by the concept of self-specialization. Each module is integrated with the base LLM, and the entire system is enhanced by a self-optimized routing mechanism. In contrast to monolithic models, which often suffer from forgetting issues when adapted or merged under fixed, static parameters, our modular design preserves the integrity and semantics of each expert. This allows for dynamic, precise handling of various target domain tasks, boosting the model’s overall capability, adaptability, and interpretability.
@@
Moreover, MiXSE clearly surpasses other strong baselines such as instance merging and weight merging, under similar settings, while offering better flexibility and interpretability. Detailed analyses highlight the critical role of the routing mechanism and the contribution of semantic experts in achieving these results. Our interpretable visualizations of routing distributions further elucidate how tasks are dynamically allocated to the most relevant experts. Lastly, we further validate that there are no issues related to forgetting unlike monolithic baselines, and that our approach can be applied to various model families and sizes.
@@ 2 Problem Statement
Here, under our problem settings, the number of parameters of $\Theta_{0}$ and $\Theta_{comp}$ should not be significantly different, necessitating that the expert modules $\Delta\Theta_{i}$ be lightweight (i.e., LoRA (Hu et al., 2022)). The available seed data are limited but can be reasonably collected (e.g., 100). Importantly, we do not assume the availability of larger/teacher models at one’s hand; instead, we aim to develop a method that enables self-improvement and is designed to be universally applicable.
@@ 3.1 Building Expert Modules through Self-Specialization
The first step of Self-MoE is creating specialized modules $\{\Delta\Theta_{i}\}_{i=1}^{n}$ for each target expertise, while adhering to the desiderata discussed in Section 2. That is, the modules should be lightweight and self-improving. We employ the concept of self-specialization (Kang et al., 2023a) where a base LLM is aligned with self-generated synthetic data for target specialization, resulting in lightweight LoRA (Hu et al., 2022) experts.
@@ 3.2 Mixture of Self-Specialized Experts
$L(\theta_{r})=$
$-\mathbb{E}_{({\text{{inst}}},\ {\text{{resp}}})\sim D}[logP_{\Theta_{0}}({\text{{resp}}}\ |\ {\text{{inst}}};\ \theta_{r},\{\Delta\Theta_{i}\}_{i\text{=1}}^{n})]$
@@
Method Knowledge Reasoning Math Coding Avg. (MMLU) (BBH) (GSM8K) (HumanEval) Base LLM 58.4 56.1 42.5 34.1 47.8 Specialized LLM for Each Capabiility Knowledge Self-Spec. 64.0 41.7 40.5 28.0 43.6 Reasoning Self-Spec. 60.1 60.2 41.0 28.7 47.5 Math Self-Spec. 59.3 58.9 50.0 36.0 51.1 Coding Self-Spec. 57.2 57.2 46.0 37.2 49.4 Merging Methods Instance Merging 62.6 57.6 53.5 36.0 52.4 TIES Merging 63.7 56.3 38.5 32.9 47.9 DARE Merging 37.7 59.6 45.0 34.8 44.3 MiXSE (Ours) 65.6  $\uparrow$ 7.2 61.1  $\uparrow$ 5.0 52.5  $\uparrow$ 10.0 37.8  $\uparrow$ 3.7 54.3  $\uparrow$ 6.5
Table 1: Main results. All models are built upon the same base LLM, Gemma-7B, and take self-improving approaches. Corresponding aligned performances of self-specialization are underscored. Each column’s best performance is highlighted in bold, while the gains achieved by our MiXSE over the base LLM are indicated.
@@ 4 Experiments and Results
We evaluate Self-MoE across diverse domains that can be categorized as knowledge, reasoning, math, and coding. We use the following benchmark datasets:
MMLU (Massive Multitask Language Understanding, 0-shot unless otherwise stated) (Hendrycks et al., 2021a): A collection of 57 academic knowledge tasks.
BBH (BIG-Bench Hard, 3-shot) (Suzgun et al., 2022): A set of 27 challenging reasoning tasks.
GSM8K (Grade School Math 8K, 8-shot) (Cobbe et al., 2021): A diverse set of grade school math word problems.
HumanEval (0-shot) (Chen et al., 2021): A hand-written evaluation set for python programming problems.
To test generalization (Section 4.4), we additionally evaluate on MATH (4-shot) (Hendrycks et al., 2021b), MBPP (3-shot) (Austin et al., 2021), NaturalQuestions (5-shot) (Kwiatkowski et al., 2019), TriviaQA (5-shot) (Joshi et al., 2017), Hellaswag (0-shot) (Zellers et al., 2019), PIQA (0-shot) (Bisk et al., 2020), and TruthfulQA (0-shot) (Lin et al., 2022).
@@
To assess the effectiveness of Self-MoE, we compare performance against several baselines:
Four Self-Specialized Models (Kang et al., 2023a): Trained on self-generated synthetic data for individual domains.
Instance Merging (Multi-Task Tuning): Leverages the aggregated synthetic data generated by self-specialization to train a model capable of handling multiple tasks simultaneously.
@@
We also contextualize these results with computationally intensive methods reported in the literature, despite indirect comparisons: BTM (Li et al., 2022), Sparse Upcycling (Komatsuzaki et al., 2023), BTX (Sukhbaatar et al., 2024), GLAN (Li et al., 2024a), Orca (Mitra et al., 2023), and Merlinite (Sudalairaj et al., 2024) in Appendix C.1.
@@ 4.1 Main Results
In Table 1, we showcase comparative benchmark results of various approaches across four specialized domains: knowledge, reasoning, math, and coding. All baselines use self-generated synthetic data based on the same Base LLM, Gemma-7B, to ensure fair comparisons.
@@
In comparison with other static merging methods like Instance Merging, TIES, and DARE, MiXSE stands out for its superior adaptability. While they attempt to combine the strengths of different specialization areas into a single model, they lack the dynamic flexibility that MiXSE offers. Notably, simple instance merging (i.e., multi-task learning), though effective in enhancing the base LLM across domains, still falls short of achieving the superior average performance of 54.3 seen with MiXSE. This validates the advantages of dynamic expert integration in a compositional system.
@@ 4.3 Routing Analysis
Understanding how MiXSE allocates tasks to its various experts is crucial for gauging both its efficiency and interpretability. By analyzing the routing distributions across four distinct domains, we aim to clearly see whether the system matches queries to the most suitable experts. Figure 3 presents the routing distributions used to solve each benchmark, where the weights are averaged across tokens and layers within individual tasks.
@@ 4.4 Generalizability Test
Concerning the potential side effect of forgetting, we extend our testing to include domains such as world knowledge, common sense, and safety, which are rarely associated with the targets directly. Our experiments show that overall, there are rarely meaningful drops in performances when applying our Self-MoE. Only a minor drop is observed with MiXSE in TriviaQA, but this is substantially less than in the case of instance merging. This suggests our approach almost maintains existing knowledge for non-targets while significantly boosting target performances, unlike monolithic baselines.
Figure 4: Results of Self-MoE applied to other LLMs.
@@ 4.7 Scaling the Number of Experts
[…]
Table 5: Comprehensive summary of relevant models for references. Detailed discussions in Appendix C.1.
@@ 5 Related Work
To offer a broader perspective, Table 5 presents a comprehensive summary of various models that, while relevant, are not directly comparable. For further discussions and a more detailed comparison, please refer to Appendix C.1.
@@
Another relevant area is model merging, which involves the weighted averaging of multiple models to form a single, aggregated model (Wortsman et al., 2022; Matena and Raffel, 2022; Ilharco et al., 2023; Jin et al., 2023). One of the leading methods, TIES (Yadav et al., 2023) tackles conflicts and parameter inconsistencies among models. DARE (Yu et al., 2024) further reduces the redundancy of parameters. However, these methods are fundamentally static in that they operate with fixed parameters once merged, which may lead to interference, lacking the dynamic flexibility that MiXSE offers.
There exist notable recent MoE models that similarly explore the utilization of semantic experts, albeit in distinct contexts (Wu et al., 2024; Muqeeth et al., 2024; Sukhbaatar et al., 2024). MOLE relies on human-labeled data, and PHATGOOSE assumes the availability of existing expert models trained by external creators and necessitates additional training for a router on the creators’ side. BTX relies on extensive pre-training, demanding significant resources, yet it as a pre-trained model holds the potential to complement our self-training approach. Unlike MOLE and PHATGOOSE, our Self-MoE framework creates experts and a router from scratch through self-improvement, while using minimal resources, as contrasted to BTX.
@@
The pursuit of enhancing the capabilities of LLMs often revolves around an instruction-tuning scheme, which can significantly boost cross-task generalizability (Ouyang et al., 2022; Su et al., 2022; Mishra et al., 2022; Wei et al., 2022). Due to the bottlenecks of expensive annotation costs which lead to limited scalability, the self-training concept (Luo, 2022) has gained attention from the community, where LLMs are aligned with automatically self-generated synthetic instructions (Wang et al., 2023; Sun et al., 2023; Li et al., 2024b). These are distinguished from distillation techniques (Hinton et al., 2015; Kang et al., 2023b), which assume a stronger teacher model (Mitra et al., 2023; Li et al., 2024a; Sudalairaj et al., 2024), limiting their applicability.
With the growing necessity to adapt generalist models to specific domains, Kang et al. (2023a) adopts the self-training paradigm for specialization, tackling that general instruction tuning is rarely effective in expert domains. While this work serves as a foundation for enhancing specialized expertise with minimal resources, we recognize the inherent trade-offs as a monolithic structure, such as potential compromises in performance outside of the specialized domains. Unlike them, our Self-MoE achieves uncompromising multiple expertise by taking a modular approach without extensive resources and adding many parameters.

### v2

@@ Abstract
We present Self-MoE, an approach that transforms a monolithic LLM into a compositional, modular system of self-specialized experts, named MiXSE (MiXture of Self-specialized Experts). Our approach leverages self-specialization, which constructs expert modules using self-generated synthetic data, each equipping a shared base LLM with distinct domain-specific capabilities, activated via self-optimized routing. This allows for dynamic and capability-specific handling of various target tasks, enhancing overall capabilities, without extensive human-labeled data and added parameters. Our empirical results reveal that specializing LLMs may exhibit potential trade-offs in performances on non-specialized tasks. On the other hand, our Self-MoE demonstrates substantial improvements (6.5%p on average) over the base LLM across diverse benchmarks such as knowledge, reasoning, math, and coding. It also consistently outperforms other methods, including instance merging and weight merging, while offering better flexibility and interpretability by design with semantic experts and routing. Our findings highlight the critical role of modularity, the applicability of Self-MoE to multiple base LLMs, and the potential of self-improvement in achieving efficient, scalable, and adaptable systems.
@@ 1 Introduction
Figure 1: Concept of Self-MoE, illustrating the transformation from a monolithic LLM to a compositional system, MiXSE, without extensive resources and addition of significant parameters. MiXSE distinguishes itself from traditional MoEs and other models in post-training, lightweight semantic experts, and/or self-generated synthetic data. The results showcase MiXSE’s improved capabilities over the base LLM (e.g., Gemma-7B) across all domains, unlike the knowledge-specialized LLM that compromises other capabilities.
Meanwhile, the increasing demand to handle domain-specific or expert-level tasks has highlighted the need for specialization of LLMs (Cheng et al., 2024; Ling et al., 2023; Feng et al., 2024). However, effective tuning often relies on high-quality, human-annotated data, which is costly and challenging to scale (Kang et al., 2023), especially in specialized domains where expertise is scarce and valuable (Wu et al., 2023). Self-specialization (Kang et al., 2024) offers a promising alternative, aligning models with self-generated synthetic data. While this technique has proven effective in cross-task generalization within a target expert domain, we posit that it may compromise performance in areas outside the target domain.
In this paper, we explore the following question: How can we build compositional LLMs that enjoy versatile expertise, while using minimal resources? We introduce Self-MoE (Figure 1), an approach that transforms a monolithic model into a compositional (Zaharia et al., 2024) system, called MiXSE (MiXture of Self-specialized Experts). This approach differs from prior MoE work using LoRA (Hu et al., 2022), which either relies on human-labeled data (Wu et al., 2024) or assumes the existence of trained modules (Huang et al., 2023; Muqeeth et al., 2024). Instead, our Self-MoE constructs individual lightweight expert modules from scratch using synthetic data, inspired by the concept of self-specialization. Each module is integrated with a shared base LLM, and the entire system is enhanced by a self-optimized routing mechanism. In contrast to monolithic models, which often suffer from forgetting issues when adapted or merged under fixed, static parameters, our modular design preserves the integrity and semantics of each expert. This allows for dynamic, precise handling of various target domain tasks, boosting the model’s overall capability, adaptability, and interpretability.
@@
Moreover, MiXSE clearly surpasses other strong baselines such as instance merging and weight merging, under similar settings, while offering better flexibility and interpretability. Detailed analyses highlight the critical role of the routing mechanism and the contribution of semantic experts in achieving these results. Our interpretable visualizations of routing distributions further elucidate how tasks are dynamically allocated to the most relevant experts. Lastly, we further validate that there are no issues related to forgetting unlike monolithic baselines, and that our approach can be applied to various model families and sizes. In summary, our key contributions are as follows:
  • We highlight the inherent limitations of monolithic model specialization, where focusing on a specific capability often comes at the cost of degrading performance in other domains.
  • We propose Self-MoE, which allows a base, monolithic LLM to upgrade into a modular system of lightweight, self-specialized experts, without requiring extensive human supervision, compute resources, or overhead in active parameters.
  • We provide comprehensive experiments and analyses across a range of benchmarks, where Self-MoE demonstrates consistent improvements with an average of 6.5%p across domains over a base LLM, outperforming various baselines. Our ablation studies validate the impact of modularity, routing strategies, and the use of self-generated synthetic data. Moreover, our analyses explore routing distributions, forgetting issues, and the applicability of our approach to five different base LLMs.
@@ 2 Problem Statement
Here, under our problem setting, the number of parameters of $\Theta_{0}$ and $\Theta_{comp}$ should not be significantly different, necessitating that the expert modules $\Delta\Theta_{i}$ be lightweight (i.e., LoRA (Hu et al., 2022)). The available seed data are limited but can be reasonably collected (e.g., 100). Importantly, we do not assume the availability of larger/teacher models at one’s hand; instead, we aim to develop a method that enables self-improvement and is designed to be universally applicable.
@@ 3.1 Building Expert Modules through Self-Specialization
The first step of Self-MoE is creating specialized modules $\{\Delta\Theta_{i}\}_{i=1}^{n}$ for each target expertise, while adhering to the desiderata discussed in Section 2. That is, the modules should be lightweight and self-improving. We employ the concept of self-specialization (Kang et al., 2024) where a base LLM is aligned with self-generated synthetic data for target specialization, resulting in lightweight LoRA (Hu et al., 2022) experts.
@@ 3.2 Mixture of Self-Specialized Experts
$L(\theta_{r})=-\mathbb{E}_{({\text{{inst}}},\ {\text{{resp}}})\sim D}[logP_{\Theta_{0}}({\text{{resp}}}\ |\ {\text{{inst}}};\ \theta_{r},\{\Delta\Theta_{i}\}_{i\text{=1}}^{n})]$
@@ 4 Experiments and Results
We evaluate Self-MoE across diverse domains categorized into knowledge, reasoning, math, and coding: MMLU (0- & 5-shot) (Hendrycks et al., 2021a), BBH (3-shot) (Suzgun et al., 2022), GSM8K (8-shot) (Cobbe et al., 2021), and HumanEval (0-shot) (Chen et al., 2021), respectively. For MMLU, we primarily employ the 0-shot setting unless otherwise specified, based on established observations (Dettmers et al., 2023; Lin et al., 2024) that tuning yields only marginal effects in the 5-shot setting for this task. To test generalization (Section 4.4), we additionally evaluate on MATH (4-shot) (Hendrycks et al., 2021b), MBPP (3-shot) (Austin et al., 2021), NaturalQuestions (5-shot) (Kwiatkowski et al., 2019), TriviaQA (5-shot) (Joshi et al., 2017), Hellaswag (0-shot) (Zellers et al., 2019), PIQA (0-shot) (Bisk et al., 2020), and TruthfulQA (0-shot) (Lin et al., 2022).
@@
To assess the effectiveness of Self-MoE, we compare performance against several baselines that use the same number of active parameters during inference:
  • Four Self-Specialized Models (Kang et al., 2024): Trained on self-generated synthetic data for individual domains: knowledge, reasoning, math, and coding.
  • Instance Merging (Multi-Task Tuning) (Chung et al., 2024): Leverages the aggregated synthetic data generated by self-specialization to train a model capable of handling multiple tasks.
@@
We also contextualize these results with computationally intensive methods reported in the literature, despite indirect comparisons: BTM (Li et al., 2022), Sparse Upcycling (Komatsuzaki et al., 2023), BTX (Sukhbaatar et al., 2024), GLAN (Li et al., 2024a), Orca (Mitra et al., 2023), and Merlinite (Sudalairaj et al., 2024) in Appendix D.1.
@@
Table 1: Main results. All models are built upon the same base LLM, Gemma-7B, taking self-improving approaches and having the same active parameters during inference. Corresponding aligned performances of self-specialization are underscored. Each column’s best performance is highlighted in bold, while the gains achieved by our MiXSE over the base LLM are indicated.
Method Active Knowledge Reasoning Math Coding Avg. Params (MMLU) (BBH) (GSM8K) (HumanEval) Base LLM 7B 58.4 56.1 42.5 34.1 47.8 Specialized LLM for Each Capabiility Knowledge Self-Spec. 7B + 0.3% 64.0 41.7 40.5 28.0 43.6 Reasoning Self-Spec. 7B + 0.3% 60.1 60.2 41.0 28.7 47.5 Math Self-Spec. 7B + 0.3% 59.3 58.9 50.0 36.0 51.1 Coding Self-Spec. 7B + 0.3% 57.2 57.2 46.0 37.2 49.4 Merging Methods Instance Merging 7B + 0.3% 62.6 57.6 53.5 36.0 52.4 TIES Merging 7B + 0.3% 63.7 56.3 38.5 32.9 47.9 DARE Merging 7B + 0.3% 37.7 59.6 45.0 34.8 44.3 MiXSE (Ours) 7B + 0.3% 65.6  $\uparrow$ 7.2 61.1  $\uparrow$ 5.0 52.5  $\uparrow$ 10.0 37.8  $\uparrow$ 3.7 54.3  $\uparrow$ 6.5
@@ 4.1 Main Results
In Table 1, we showcase comparative benchmark results of various approaches across four specialized domains: knowledge, reasoning, math, and coding. All baselines use self-generated synthetic data based on the same Base LLM, Gemma-7B, and LoRA for tuning to ensure fair comparisons.
@@
In comparison with other static merging methods like Instance Merging, TIES, and DARE, MiXSE stands out for its superior adaptability. While they attempt to combine the strengths of different specialization areas into a single model, they lack the dynamic flexibility that MiXSE offers. Notably, simple instance merging (i.e., multi-task tuning), though effective in enhancing the base LLM across domains, still falls short of achieving the superior average performance of 54.3 seen with MiXSE. This validates the advantages of dynamic expert integration in a compositional system.
@@ 4.3 Routing Analysis
Understanding how MiXSE allocates tasks to its various experts is crucial for gauging its interpretability. By analyzing the routing distributions across four distinct domains, we aim to see whether the system matches queries to the most suitable experts. Figure 3 presents the routing distributions used to solve each benchmark, where the weights are averaged across tokens and layers within individual tasks.
@@ 4.4 Generalizability Test
Concerning the potential side effect of forgetting, we extend our testing to include domains such as world knowledge, common sense, and safety, which are rarely associated with the targets directly. Our experiments show that overall, there are rarely meaningful performance drops when applying our Self-MoE. Only a minor drop is observed with MiXSE in TriviaQA, but this is substantially less than in the case of instance merging. This suggests our approach almost maintains existing knowledge for non-targets while significantly boosting target performances, unlike monolithic baselines.
Figure 4: Results of Self-MoE w/ other LLMs.
@@ 4.7 Scaling the Number of Experts
4.8 Discussion on the Overhead of Self-MoE
One possible concern in adapting LLMs into compositional systems using Self-MoE is the potential introduction of overhead. Here, we discuss this aspect in detail, emphasizing that the additional overhead of Self-MoE is minimal while yielding significant performance gains. Essentially, the expert modules in Self-MoE are lightweight LoRA modules, contributing only about 1% additional parameters (total) for four experts, as detailed in Table 5 (Total Params). These experts are sparsely activated, which maintains low active parameters (7B + 0.3%) during inference, thus efficiently minimizing inference overhead. In contrast, traditional MoE models like Mixtral (Jiang et al., 2024) and BTX (Sukhbaatar et al., 2024) typically employ a feedforward network (FFN) layer for each expert, resulting in a significant proportional increase in total parameters as the number of experts grows, as indicated in Table 5, which demands much more memory for model loading. The design choice in Self-MoE leads to better scalability and resource efficiency, especially when the number of experts is scaled to incorporate numerous domains of expertise.
@@ 5 Related Work
To offer a broader perspective, Table 5 presents a comprehensive summary of various models that, while relevant, are not directly comparable. For further discussions and a more detailed comparison, please refer to Appendix D.1.
@@
Another relevant area is merging, involving the weighted averaging of multiple models to form a single, aggregated model (Wortsman et al., 2022; Matena & Raffel, 2022; Ilharco et al., 2023; Jin et al., 2023). One of the leading methods, TIES (Yadav et al., 2023) tackles conflicts and parameter inconsistencies among models. DARE (Yu et al., 2024) further reduces the redundancy of parameters. However, these methods are fundamentally static in that they operate with fixed parameters once merged, which may lead to interference, lacking the dynamic flexibility that MiXSE offers.
There exist notable recent MoE models that similarly explore the utilization of semantic experts, albeit in distinct contexts (Gururangan et al., 2022; Wu et al., 2024; Muqeeth et al., 2024; Sukhbaatar et al., 2024). MOLE relies on human-labeled data, and PHATGOOSE assumes the availability of existing expert models trained by external creators and necessitates additional training for a router on the creators’ side. DEMix and BTX rely on extensive pre-training, demanding significant resources, yet it as a pre-trained model holds the potential to complement our self-training approach. Unlike MOLE and PHATGOOSE, our Self-MoE framework creates experts and a router from scratch through self-improvement, while using minimal resources, as contrasted to DEMix and BTX.
Table 5: Comprehensive summary of relevant models for references. Detailed discussions are provided in Appendix D.1.
[…]
@@
The pursuit of enhancing the capabilities of LLMs often revolves around an instruction-tuning scheme, which can significantly boost cross-task generalizability (Ouyang et al., 2022; Su et al., 2022; Mishra et al., 2022; Wei et al., 2022). Due to the bottlenecks of expensive annotation costs which lead to limited scalability, the self-training concept (Luo, 2022) has gained attention from the community, where LLMs are aligned with automatically self-generated synthetic instructions (Wang et al., 2023; Sun et al., 2023; Li et al., 2024b). These are distinguished from distillation techniques (Hinton et al., 2015; Kang et al., 2023), which assume a stronger teacher model (Mitra et al., 2023; Li et al., 2024a; Sudalairaj et al., 2024), limiting their applicability.
With the growing need to adapt generalist models to specific domains, Kang et al. (2024) adopts the self-training for specialization, tackling that general instruction tuning is rarely effective in expert domains. While this work lays a foundation for enhancing specialized expertise with minimal resources, we recognize inherent trade-offs in a monolithic structure, such as performance compromises outside specialized domains. Conversely, our Self-MoE achieves uncompromising multiple expertise with a modular approach without extensive resources and adding many parameters.